OpenStudy (anonymous):

16*8+10/2= please help dnt understand

OpenStudy (anonymous):

16*8 = 128 --> 128 + 10/2 = 128 + 5 = 133

OpenStudy (stormfire1):

This is a question testing your knowledge of the order of operations (PEMDAS - Parenthesis, Exponents, Multiplication, Division, Addition and Subtraction...in that order). Since you only have multiplication and addition in the problem, using the order of operations above means that multiplication comes first, left-to-right. Multiple 16 * 8...then divide 10/2...THEN you can add them together to get the answer (133)

OpenStudy (anonymous):

thankx soooo much i will definitelytry harder...lol...